# Singular parametric oscillators from the one-parameter Darboux transformation of the classical harmonic oscillator

### v1

@@ Abstract
The singular parametric oscillators obtained from the one-parameter Darboux deformation/transformation effected upon the classical harmonic oscillator are introduced and discussed in some detail using $\sin(\omega_{0}t)$ and $\cos(\omega_{0}t)$ as seed solutions. The corresponding Ermakov-Lewis integrability problem of these parametric oscillators is also studied. It is shown that the Ermakov-Lewis invariants do not depend on the deformation parameter.
@@ I Introduction
One-parameter Darboux transformations of homogeneous second-order linear differential equations have been an interesting topic in the literature since four decades now. Different from the common Darboux transformations which are based on a particular Riccati solution, they are based on the general Riccati solution and therefore generate whole families of Darboux covariant equations. In a quantum mechanical context, this covariance has been used to generate many exactly solvable Schrödinger problems that are isospectrally related to a given exactly solvable quantum eigenvalue problem with the parameter related to the changed character of the boundary conditions. This kind of parametric Darboux transformations have been applied in many areas of physics pM ; pF ; psp ; rosu96-1 ; rosu96-2 ; rosu97 ; r2000 ; RMC including the free damped harmonic oscillator in classical mechanics rr98 ; re01 , but, to the best of our knowledge, the pure harmonic oscillator has not been studied in this approach.
In this short paper, our aim is to provide a detailed study of the parametric oscillators obtained by applying the one-parameter Darboux transformation to the classical harmonic oscillator,
@@
as discussed in the one-parameter supersymmetric framework in the recent work rm23 . In (1.2), $\tilde{\varphi}\approx 0.4812$ is the natural logarithm of the golden ratio $\varphi\equiv\frac{1+\sqrt{5}}{2}\approx 1.6180$ and the subindex refers to the odd and even generations of Fibonacci numbers. In rm23 , the one-parameter supersymmetric approach of (1.2) has been motivated by the idea of introducing parametric Darboux-deformed Fibonacci numbers due to the fact that the hyperbolic function solutions are the odd and even Binet expressions of the Fibonacci numbers as previously pointed out by Faraoni and Atieh FA .
The rest of the present paper is structured as follows. Section II contains a brief review of the parametric supersymmetric approach. The application to equation (1.1) is described in section III. The associated Ermakov problem and the corresponding Ermakov-Lewis invariants are discussed in section IV. Finally, the conclusions end up the paper.
@@ II Brief review of the one-parameter Darboux transformation
$\left(D^{2}+f(t)\right)F=0~{},\qquad D=\frac{d^{2}}{dt^{2}}~{},\qquad f(t)=\omega_{0}^{2}~{},$ (2.1)
@@ III The one-parameter Darboux transformation of the classical harmonic oscillator
Figure 1: The square of the odd Darboux-deformed angular frequency parameter (up) and the odd deformed linearly independent solutions (down) for $\omega_{0}=1$, and $\gamma=0.5$ and $1$
@@
where again the second solution has been obtained from the reduction of order formula with the Wronskian $W=\omega_{0}$ corresponding to the undeformed pair ordered as $\{\cos(\omega_{0}t),\sin(\omega_{0}t)\}$.
@@ IV The associated Ermakov-Pinney equations and the Ermakov-Lewis invariants
In the odd case, we take $u_{1}={\cal F}_{o,1}(\omega_{0}t;\gamma)$ and $u_{2}={\cal F}_{o,2}(\omega_{0}t;\gamma)$ from (3.3) and using (4.3), we have
@@
Plots of the Ermakov-Pinney solutions ${\cal V}_{o}$ and ${\cal V}_{e}$ for $\gamma=0.5$ and $\gamma=1$ are displayed in Fig. 6.
@@
V Conclusions
In this paper, we have introduced and discussed in some detail an interesting class of singular parametric oscillators obtained through the one-parameter Darboux deformation/transformation of the classical harmonic oscillator. In an innovative way, the oscillatory singular solutions have been written as amplitude-modulated solutions of the sine and cosine functions that have been used as seed solutions. Moreover, for the corresponding Ermakov-Pinney equations, the solutions have been also written in the amplitude-modulated form making easier the calculation of the Ermakov-Lewis invariants. The latter invariants have been shown to depend on the superposition constants of the linearly independent harmonic solutions, but not on the deformation parameter.
@@ Acknowledgments
The second author acknowledges the financial support of CONAHCyT through a doctoral fellowship.

### v2

# Ann. Phys. 470 (2024) 169830

@@ Abstract
The singular parametric oscillators obtained from the one-parameter Darboux deformation/transformation effected upon the classical harmonic oscillator are introduced and discussed in some detail using $\sin(\omega_{0}t)$ and $\cos(\omega_{0}t)$ as seed solutions. The corresponding Ermakov-Lewis integrability problem of these parametric oscillators is also studied. It is shown that their Ermakov-Lewis invariants do not depend on the deformation parameter and are singularity-free.
@@ I Introduction
One-parameter Darboux transformations of homogeneous second-order linear differential equations have been an interesting topic in the literature since four decades now. Different from the common Darboux transformations which are based on a particular Riccati solution, they are based on the general Riccati solution and therefore generate whole families of Darboux covariant equations. In a quantum mechanical context, this covariance has been used to generate many exactly solvable Schrödinger problems that are isospectrally related to a given exactly solvable quantum eigenvalue problem with the parameter related to the changed character of the boundary conditions. This kind of parametric Darboux transformations has been applied in many areas of physics pM ; pF ; psp ; cks ; rosu96-1 ; rosu96-2 ; rosu97 ; r2000 ; RMC ; KKYK including the free damped harmonic oscillator in classical mechanics rr98 ; re01 , but, to the best of our knowledge, the pure harmonic oscillator has not been studied in this approach. Mathematically, the parametric Darboux transformations turn these simple autonomous equations into nonautonomous equations pertaining to the vast area of time-dependent oscillators of ever lasting research interest td1 ; td2 ; td3 . On the other hand, at the formalism level, supersymmetric classical Hamiltonians and Lagrangians have been studied in the literature employing additional Grassmannian variables bagchibook ; bissoni86 ; ioffe06 .
In this paper, our aim is to provide a detailed study of the parametric oscillators obtained by applying the one-parameter Darboux transformation to the classical harmonic oscillator,
@@
as discussed in the one-parameter supersymmetric framework in the recent work rm23 . In (1.2), $\tilde{\varphi}\approx 0.4812$ is the natural logarithm of the golden ratio $\varphi\equiv(1+\sqrt{5})/2\approx 1.6180$ and the subindex refers to the odd and even generations of Fibonacci numbers. In rm23 , the study of the one-parameter Darboux deformation of (1.2) has been motivated by the idea of introducing parametric Darboux-deformed Fibonacci numbers by taking into account that the hyperbolic function solutions are the odd and even Binet expressions of the Fibonacci numbers as previously pointed out by Faraoni and Atieh FA .
The rest of the paper is structured as follows. Section II contains a brief review of the parametric supersymmetric approach. The application to equation (1.1) is described in section III. The associated Ermakov problem and the corresponding Ermakov-Lewis invariants are discussed in section IV. Finally, some discussions and conclusions end up the paper.
@@ II Brief review of the one-parameter Darboux transformation
$\left(D^{2}+f(t)\right)F=0~{},\qquad D^{2}=\frac{d^{2}}{dt^{2}}~{},\qquad f(t)=\omega_{0}^{2}~{},$ (2.1)
@@ III The one-parameter Darboux transformation of the classical harmonic oscillator
Figure 1: The square of the odd Darboux-deformed angular frequency parameter (up) and the odd deformed linearly independent solutions (down) for $\omega_{0}=1$, and $\gamma=0.5$ and $1$. (In all plots of the paper, we set $\omega_{0}=1$).
@@
where again the second solution has been obtained from the reduction of order formula with the Wronskian $W=\omega_{0}$ corresponding to the non-deformed pair ordered as $\{\cos(\omega_{0}t),\sin(\omega_{0}t)\}$.
@@ IV The associated Ermakov-Pinney equations and the Ermakov-Lewis invariants
In the odd case, we take $u_{1}={\cal F}_{o,1}(\omega_{0}t;\gamma)$ and $u_{2}={\cal F}_{o,2}(\omega_{0}t;\gamma)$ from (LABEL:3.17) and using (4.3), we have
@@
Plots of the parametric Ermakov-Pinney solutions ${\cal V}_{o}$ and ${\cal V}_{e}$ for $\gamma=1$ and $\gamma=2$ are displayed in Fig. 6.
@@
Another interesting feature of the Ermakov-Lewis invariant for the Darboux-deformed parametric oscillators is that it is free of both trigonometric and parameter singularities. This is due to the special form of the deformed solutions and the corresponding Pinney solutions which provide opposite oscillations around the constant value of the invariant, see Figs. 8 and 9.
To this end, following ray1979 , we discuss the adiabatic regime of these singular parametric oscillators. This dynamical regime, defined by the parameter $|\theta|=|d{\omega/dt}|/\omega^{2}(t)\ll 1$, is an important technological characteristic of parametric oscillators and is controlled by its own quasi-invariant known as the adiabatic invariant
$I_{{\rm ad}}=\frac{E(t)}{\omega(t)}~{},$ (4.14)
where the energy of the oscillator is given by
$E(t)=\frac{1}{2}\big{[}(d{\cal F}/dt)^{2}+\omega^{2}(t){\cal F}^{2}\big{]}~{}.$ (4.15)
Lewis discovered that the adiabatic ‘invariant’ is the leading term in the adiabatic series expansion of the Ermakov-Lewis invariant lewis1968 . However, an adiabatic regime for the parametric Darboux-deformed modes can exist only for the modes ${\cal F}_{o1}$ and ${\cal F}_{e1}$ since their energy is decreasing in time, but the adiabaticity condition on the parameter $\theta$ is better satisfied only at high values of the parameter $\gamma$, see Fig. 10, where the plots show that $|\theta|<2$% for $\gamma>$ 100 in both cases.
@@
Figure 8: The oscillations of the two terms in the Ermakov-Lewis invariants in the first case from the previous figure. The vertical dashed lines indicate the time location of the parameter-induced singularities in the deformed functions and the corresponding Pinney functions at $t=-1.78$ and $t=-4.34$ for $\gamma=1$ and $\gamma=2$, respectively.
Figure 9: The oscillations of the two terms in the Ermakov-Lewis invariants in the even case from Fig. 7. The vertical dashed lines indicate the time location of the parameter-induced singularities in the deformed functions and the corresponding Pinney functions at $t=-2.484$ and $t=-3.602$ for $\gamma=1$ and $\gamma=2$, respectively.
Figure 10: The adiabaticity parameter $|\theta|$ for the one-parameter Darboux-deformed modes ${\cal F}_{o1}$ and ${\cal F}_{e1}$ for various values of $\gamma$.
V Discussion and Conclusions
In this paper, we have introduced and discussed in some detail an interesting class of singular parametric oscillators obtained through the one-parameter Darboux deformation/transformation of the classical harmonic oscillator. In an innovative way, the oscillatory singular solutions have been written as amplitude-modulated solutions of the sine and cosine functions that have been used as seed solutions. Moreover, for the corresponding Ermakov-Pinney equations, the solutions have been also written in the amplitude-modulated form making easier the calculation of the Ermakov-Lewis invariants. The latter invariants have been shown to depend on the superposition constants of the linearly independent harmonic solutions, but not on the deformation parameter. Significant insight has been accomplished by showing that the Ermakov-Lewis invariants remain true dynamical invariants for these Darboux-deformed parametric oscillators despite their parametric and trigonometric singularities. This is due to the opposite functional behaviours of the quotient term and the Wronskian-like term in the formula of the Ermakov-Lewis invariant that lead to cancelations of the singularities. An adiabatic regime may be considered, but only for the decaying modes at high values of the parameter $\gamma$.
The more general case of this kind of parametric oscillators possessing constant friction or gain can be easily analysed by adding a constant shift to the Riccati solution, a procedure that we described in rk2011 , or by direct factorization as in rr98 , which is similar to the nonlinear factorizations of Rosu and Cornejo-Pérez rcp1 ; rcp2 , which can be also adapted to differential equations with nonautonomous friction-gain terms.
One obvious application of the results of this paper, including those related to the Ermakov-Lewis approach, is in cosmology, where the Riccati/supersymmetry approach in the conformal time variable has been long considered in the FRW barotropic models r2000 ; r98 ; r02 ; rk2013 and isotropic scalar field cosmologies hlm2014 . The even case considered in the present paper corresponds to the evolution of a closed vacuum-filled FRW universe in conformal time. Regarding the usage of the Ermakov-Lewis invariants in this area, we mention the comment paper of Ray ray1979 , who related the cosmological particle production in the adiabatic regime to the Ermakov-Lewis invariant of parametric oscillator modes with the angular frequency depending on the time-dependent metric of cosmological models with an initial singularity. We also recall here the recent surge of interest in this approach in a number of stimulating, cosmology-sided papers bkpt24 ; crs24 ; barb06 , but with some results that can be adapted to the quantum-to-classical transition at microscopic scales.
@@ Acknowledgments
The authors thank the referee for inspiring remarks that help them to greatly improve the quality of this research. The second author acknowledges the financial support of CONAHCyT-Mexico through a doctoral fellowship at the early stage of this work.